# The Generation Gap: Exploring Age Bias in Large Language Models

### v1

@@ Abstract
In this paper, we explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories. Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics. Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts. Our findings highlight the age bias in LLMs and provide insights for future work.
The Generation Gap: Exploring Age Bias in Large Language Models
Siyang Liu1   Trish Maturi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
@@ 1 Introduction
Figure 1: Age-related bias in LLMs on thirteen human value categories. Human values in this figure refer in particular to the US groups. Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM values and human values as age increases. A positive trend coefficient signifies the widening gap observed from younger to older age groups, thus indicating a model leaning towards younger age groups.
@@
We advocate for a heightened awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values. We also emphasize the complexities involved in calibrating prompt engineering to effectively address this bias.
2 Social Biases in LLMs
Recent advancements in LLMs have led to human-level or even surpassing performance across various NLP tasks Brown et al. (2020); Radford et al. (2019); Ouyang et al. (2022). However, there is a growing concern regarding the presence of social bias in these models Kasneci et al. (2023), especially as certain biases could result in discrimination and emotional harm to the impacted users. Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals, while struggling with predictive capabilities concerning Black and Asian groups Sun et al. (2023). Additionally, recent studies by Santurkar et al.; McGee; Atari et al. consistently highlight a left-leaning or democratic political inclination among LLM models. Despite extensive scrutiny on LLM bias Santurkar et al. (2023); Sun et al. (2023); Duan et al. (2024), the age-related preferences of LLMs remain less explored, and call for a comprehensive investigation for a more equitable and inclusive technological landscape.
Figure 2: Alignment rank of values of ChatGPT over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with ChatGPT in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
@@ 3 Analytic Method
3.1 Dataset Processing
We derive human values across age demographics and create prompts utilizing data from the 7th wave of the World Values Survey (WVS)  [^1] Statisitcs of WVS is accessible via Document-Online analysis. Haerpfer et al. (2020). The survey systematically probes 94278 individuals globally on thirteen categories, covering a range of social, political, economic, religious, and cultural values. To assess human values, we group the respondents by age group  [^2] Age groups are recorded as 18-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country. Subsequently, we compute the average values for each age group and country to represent their respective cohorts.
@@ 3.2 Prompting
We conduct our analysis on four LLMs: ChatGPT (GPT-3.5-turbo 0613), InstructGPT (GPT-3.5-turbo-instruct) OpenAI (2023), FLAN-T5XXL Chung et al. (2022), and FLAN-UL2 Tay (2023), spanning both open-source and close-source, chat-based and completion-based, as well as decoder-only and encoder-decoder architecture LLMs.
@@
We identify three key components for each inquiry in the survey: context, question ID&content, and options. In addition, we add format variation—also known as spurious variation—into prompts, as previous research Shu et al. (2023) suggests inconsistent performance in LLMs after receiving a minor prompt variation.
By altering both the format and the order of inquiry components, we build a set of eight distinct prompt per inquiry as shown in Tab 3. We utilize the average outcomes across multiple prompt variations for our analysis to make our probing method more robust. Due to the variety in LLM’s capability in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 1. We observe a high agreement among responses induced by different prompts. Specifically, though we observed LLM’s instability to prompt variations (See analysis in Appx B), for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options.
Unexpected Reply Type Example Coping Method
returning null value { ”Q1”: null} map null into missing code -2
redundant texts ”Answer = {‘Q1’, 1}” extract the json result
substandard json Q1:‘1’ manually correct
inconsistent redundancy {‘Q1’:1} {‘Q1’:2} pick the firstly-shown item
@@ 3.3 Measures
We represent values belonging to a certain category as a vector. Each question in the WVS questionnaire is treated as a dimension (See the number of questions for each category in Tab 2). Calculating the distance between two vectors upon the original dimensionality often leads to a sensitivity to the dimensions of outliers or trivial information. Instead, we utilize a principle component analysis (PCA) Tipping and Bishop (1999) on a range of value vectors for learning smooth representations. Specifically, we collect 372 different value vectors representing people across 62 countries and six age groups. After performing min-max normalization and normal standardization, we conduct a PCA to reduce the vectors to a dimensionality of two. Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value vector represented by the $i$th age group be $x_{i},y_{i}]$. We derive three metrics below for our further analysis:
Euclidean Distance, the distance between two value vectors.
$d_{i}=\sqrt{{(x_{LLM}-x_{i})}^{2}+{(y_{LLM}-y_{i})}^{2}},$
where $(x_{LLM},y_{LLM})$ represents values of LLM. Alignment Rank, the reverse rank of distance between LLM values and people across six age groups.
$r_{i}=argsort(argsort([d_{1},...,d_{6}]))[i]$
Trend Coeficient, the slope of the gap between LLM values with human across six age groups.
$r_{i}=\beta+\alpha i$
$\alpha=\arg\max_{\alpha}(\sum_{i=1}^{6}(r_{i}-(\beta+\alpha i))^{2})$
@@ 4 Aligning with Which Age on Which Values?
As shown in Fig 1, we observe a general inclination of four popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient. Fig 2 exemplifies this bias for ChatGPT, where the model tends to have a higher alignment rank on younger groups on the most categories, despite few exceptions (see the left case in Fig 3). The results on other LLMs and other countries can be found in Appx D and E.
@@ 5 The Effect of Adding Identity in Prompts
We also showcase a successful calibration example for a question from the political interest WVS section in Fig 5. The value pyramid illustrates LLMs’ responses for different age ranges compares to the answers from the U.S. population. When age is factored into the LLM prompt, the LLM’s views are more aligned with the U.S. population of that respective age group, as it reports higher frequency using radio news for the older group.
@@ Appendix A World Value Survey
The WVS [^3] https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values. See the statistics of inquiries in Fig 2. Note that we remove ten of them that requires demographic information, as these are impossible for applying to an LLM lacking demographic data, and keep 249 inquiries as our final choices for prompting.
@@
(b) Prompt Orders
@@ Appendix B The Instability of LLM Outputs Due to Prompt Variations
Appendix C Prompting Details
@@
The cost of API calling from Closed-coursed LLMs is less than 5 dollars. For the deployment of FLAN-T5-XXL and FLAN-UL2 models, we ran either model on a single A40 GPU with float16 precision. When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 43.
Appendix D Results on Other LLMs
@@
Appendix E Results on Other Countries
@@
Table 4: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
Table 5: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
Table 6: Alignment rank of LLMs over different age groups in Gemany and Great Britain. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
Table 7: Alignment rank of LLMs over different age groups in Indonesia and Malaysia. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT

### v2

# The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models

@@ Abstract
In this paper, we explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories. Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics, especially in the US. Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts. Our findings highlight the age bias in LLMs and provide insights for future work. Materials for our analysis will be available via anonymous.github.com
The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models
Siyang Liu1   Trish Maturi1   Bowen Yi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
@@ 1 Introduction
Figure 1: Age-related bias in LLMs on thirteen human value categories. Human values in this figure refer in particular to the US groups. Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM and human values as age increases. A positive trend coefficient signifies the widening gap observed from younger to older groups, thus indicating a model leaning towards younger age groups. Significant test is detailed in Appx G
@@
We advocate for increased awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values. We also emphasize the complexities involved in calibrating prompts to effectively address this bias.
2 Related Work
Due to the recent advancements in LLMs in manifesting human-level performance across various tasks Brown et al. (2020); Radford et al. (2019); Ouyang et al. (2022), there is a growing concern regarding the presence of social bias in these models Kasneci et al. (2023). Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals Sun et al. (2023), and political inclination Santurkar et al. (2023); McGee (2023); Atari et al. (2023). Despite extensive scrutiny on LLM bias Santurkar et al. (2023); Sun et al. (2023), the age-related preferences of LLMs remain less explored. Previous work has mentioned age as one of multi-facets of bias in LLM performance Kamruzzaman et al. (2023); Haller et al. (2023); Draxler et al. (2023); Levy et al. (2024); Oketunji et al. (2023), while lacking a direct study on the age aspect. Recent research Duan et al. (2024) publishes an evaluation for well-known LLMs on age bias through 50 multi-choice questions; unlike it focuses on discriminatory narratives towards specific age groups, our investigation is running at an implicit level. We argue that understanding the underlying value systems is crucial, as the value discrepancies between users and LLMs can significantly impact their adoption of LLMs, even though the explicit discrimination is rectified, as exemplified in our discussion on technology attitudes in Sec 1.
Figure 2: Alignment rank of values of LLMs over different age groups in specific Countries. See results on more models and countries in Appx E and F . Rank 1 on an specific age group represents that this age group has the narrowest gap with LLM in values. A increasing monoticity indicates a closer alignment towards younger groups.
@@ 3 Analytic Method
3.1 Human Data Acquisition
Dataset.
We derive human values utilizing the 7th wave of the World Values Survey (WVS) Haerpfer et al. (2020). The survey systematically probes 94k individuals globally on 13 categories, covering a range of social, political, economic, religious, and cultural values. See an introduction of WVS in Appx A. Each inquiry is a single-choice question. Responses are numeric, quantifying the inclination on the options, e.g., “1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree”. Negative number is possible for coding exceptions such as “I don’t know”. To assess human values, we group the respondents by age group  [^1] Age groups are recorded as 18-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country. Subsequently, we compute the average values for each age group and country to represent their respective cohorts, ignoring the invalid negative numbers.
@@ 3.2 Prompting
We conduct our analysis on six LLMs, as introduced in Tab 1.
Model (Version) Features
ChatGPT(GPT-3.5-turbo 0613)
InstructGPT (GPT-3.5-turbo-instruct)
Mistral (mistral-7B-v0.1)
Vicuna (vicuna-7b-v1.5)
FLAN-T5 (flan-t5-xxl)
FLAN-UL2 (flan-ul2)
Table 1: Model description.
: commercial models,
: open models,
: chat-based,
: completion-based,
: RLHF, and
: training with instructions.
@@
We identify three key components for each inquiry in the survey: context, question ID&content, and options. To ensure robustness, we made several format variations for the prompt [^2] Despite adopting format variations, we were cautious to not include major changes as the content and structure of WVS were carefully designed by sociologists and professionals. (e.g., alter wordings and change order of components), as previous research Shu et al. (2023); Röttger et al. (2024); Beck et al. (2023) uncovered inconsistent performance in LLMs after receiving a minor prompt variation. Eventually, we build a set of eight distinct prompts per inquiry. Please see prompt design details in Tab 3 in Appendix. Through a careful analysis on the prompt responses (Appx B), We observe unstableness of LLM’s responses to prompt variations. However, multiple prompt trials assists with achieving a convergence point. On 95.5% of questions, more than half of the eight prompts led to responses centered on the same choice or adjacent options, and thus we believe it is acceptable to consider the average of the outcomes across the eight prompt variations as the LLM’s final responses to WVS. In addition, due to the instability of LLMs in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 4. In the process of averaging responses, we ignore the invalid negative numbers, as we did in calculating human values. For reproducing our work, parameter setting and prompting details are reported in Appendix D.
@@ 3.3 Measures
We use vector $V_{c}$ to represent values belonging to a certain category $c$. Each question in the WVS questionnaire is treated as a dimension:
[…]
where $r_{i}$ is a numeric response to the $i$th question in the section of $c$, and $n_{c}$ denotes the total question number. Note the acquisition of numeric responses for human groups and LLM has been illustrated in Sec 3.1 and 3.2.
By collecting 372 value vectors that represent people across 62 countries and 6 age groups, along with a value vector for the LLM to compare, we utilize principle component analysis (PCA) Tipping and Bishop (1999) on totally 373 value vectors for representation learning. We acquire value representations for all groups with the dimensionality of three. Our consideration of using PCA is added in Appx C.
[…]
Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value representation for the $i$th age group be $x_{c,i},y_{c,i},z_{c,i}]$. We derive three metrics below for our further analyses:
Euclidean Distance, the distance between two value representations.
[…]
where $(x_{c,M},y_{c,M},z_{c,M})$ represents values of LLM on category $c$.
Alignment Rank, the ascending rank of distances between LLM values and people across six age groups.
[…]
Trend Coefficient, the slope of the value gap between LLM and humans across six age groups. $\alpha$ is the slope we would like to fit by linear regression.
[…]
@@ 4 Aligning with Which Age on Which Values?
As shown in Fig 1, we observe a general inclination of popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient. Fig 2 exemplifies the bias for LMMs across six age groups in several countries. Due to the limited paper pages, results on other LLMs and countries can be found in Appx E and F. Significant testing procedure is available in Appx G. We observe that in the US and China, as countries of large population, the models tend to have a higher alignment rank on younger groups on the most categories, despite few exceptions (e.g., happiness and well-being). However, in Ethiopia and Nigeria (Tab 8), the inclination is less evident. We leave this phenomenon for future study.
@@ 5 The Effect of Adding Identity in Prompts
We also showcase a successful calibration example for a question about the source of acquiring information in Fig 5. The value pyramid illustrates LLMs’ responses for different age ranges compares to the answers from the U.S. population. When age is factored into the LLM prompt, the LLM’s views are more aligned with the U.S. population of that respective age group, as it reports higher frequency using radio news for the older group.
@@ Appendix A World Value Survey
The WVS [^3] https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values. We share a page of WVS questionnaire in Tab 6. See the statistics of inquiries in Fig 2. Demographic statistics of WVS is accessible via Document-Online analysis. Note that we remove ten of them that requires demographic information, as these are impossible for applying to an LLM lacking demographic data, and keep 249 inquiries as our final choices for prompting.
@@
(b) Eight Prompts with Changing Orders
An Example Prompt for Order 1 2.2 3.1 4.1
For each of the following statements I read out, can you tell me how strongly you agree or disagree with each. Do you strongly agree, agree, disagree, or strongly disagree? Q1:One of my main goals in life has been to make my parents proud. Options: 1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree. Answer in JSON format, where the key should be a string of the question id (e.g., Q1), and the value should be an integer of the answer id.
(c) Example Prompt
@@
Unexpected Reply Type Example Coping Method
returning null value { ”Q1”: null} map null into missing code -2
redundant texts ”Answer = {‘Q1’, 1}” extract the json result
substandard json Q1:‘1’ manually correct
inconsistent redundancy {‘Q1’:1} {‘Q1’:2} pick the firstly-shown item
@@ Appendix B The Instability of LLM Outputs Due to Prompt Variations
Appendix C Reasons of Applying PCA
  1. 1.
    Each question in WVS ought not to be equally important. Furthermore, for the questions belonging to a certain category, they correlate with each other. To this end, we need to find out the principle components among multi inquiries.
  2. 2.
    PCA here is also used as an unsupervised representation learning method. Compared to utilizing original data, the representations learned from hundreds of comparable examples (372 value vectors from different country and age groups) will mitigate the curse of dimensionality and other undesired properties of high-dimensional spaces. Other representation learning methods are also applicable. As the medium number of original dimensionality for all categories is 11, PCA is enough to handle the learning problem.
Appendix D Prompting Details
@@
The cost of API calling from Closed-coursed LLMs is less than 5 dollars. For the deployment of open-sourced models, we ran either model on a single A40 GPU with float16 precision. When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 42.
Appendix E Results on Other LLMs
@@
Appendix F Results on Other Countries
@@
Appendix G Significant Test
In this section, we conduct two kinds of significant tests to support our study: (1) we use MANOVA to test the significant difference among human values from different age groups, and (2) we use t-distribution to test the significant tendency of LLMs towards younger group. Notes our focus lies in characterizing the inclination of LLM values towards specific age groups. So, we are claiming significant tendency instead of claiming LLMs significantly assemble of any specific age groups.
G.1 Significant Test for the Discrepancy among Human Age Groups
Our analysis should base on a reasonable precondition that in WVS, human values significantly diverse cross different age groups. We used MANOVA (multivariate analysis of variance) to test the significant difference of human values cross all age groups, as shown below:
Null hypothesis ($H_{0}$): the age group has no effect on any responses to the survey questions
Statistics: Wilks’ lambda
Result: See Tab 5. In conclusion: We reject the null hypothesis with p-value ¡ 1e-4
Country Value Num DF Den DF F Value Pr ¿ F (p-value)
US 0.07 176.00 1631.00 124.82 0.0000*
China 0.06 184.00 2068.00 164.16 0.0000*
Germany 0.05 118.00 1048.00 173.11 0.0000*
Great British 0.06 118.00 1607.00 220.91 0.0000*
Indonesia 0.09 201.00 2310.00 113.78 0.0000*
Malaysia 0.09 254.00 1022.00 42.43 0.0000*
Ethiopia 0.16 127.00 843.00 34.02 0.0000*
Nigeria 0.13 176.00 614.00 23.18 0.0000*
Table 5: P-values of value difference among different age groups in specific countries. * indicates p-value¡1e-4
G.2 Significant Test for Trend Coefficient
As it may be hard to interpret the trend coefficient in Fig 1 on some categories (e.g., perception of corruption). Despite its bias towards younger/older, it may not be a significantly meaningful number. We add significance testing for the linear regression on trend coefficient.
Null hypothesis ($H_{0}$): $\alpha=0$, where is the trend coefficient fitted by a linear regression model presented in Sec 3.3.
Statistics: t distribution.
Results: see Tab 6.
Category ChatGPT InstructGPT Mistral Vicuna Flan-t5 Flan-ul
Social Values, Norm, Stereotypes 0.33 0.111 0.208 0.072* 0.005* 0.042*
Happiness and Wellbeing 0.042* 0.208 0.005* 0.005* 0.005* 0.005*
Social Capital, Trust and Organizational 0.397 0.872 0.005* 0.000* 0.042* 0.397
Economic Values 0.000* 0.468 0.872 0.468 0.623 0.042*
Perceptions of Corruption 0.704 0.072* 0.019* 0.072* 0.019* 0.005*
Perceptions of Migration 0.072* 0.042* 0.005* 0.266 0.000* 0.156
Perceptions of Security 0.042* 0.000* 0.000* 0.000* 0.000* 0.000*
Index of Postmaterialism 0.623 0.787 0.397 0.111 0.787 0.005*
Perceptions about Science and Technology 0.329 0.468 0.329 0.005* 0.329 0.623
Religious Values 0.111 0.544 0.005* 0.005* 0.005* 0.019*
Ethical Values 0.000* 0.000* 0.000* 0.000* 0.072* 0.000*
Political Interest and Political Participation 0.208 0.872 0.000* 0.000* 0.208 0.329
Political Culture and Political Regimes 0.000* 0.000* 0.000* 0.005* 0.957 0.872
Table 6: P-values of trend coefficients for each model on each value category. * indicates p-value¡0.1
@@ G.2 Significant Test for Trend Coefficient
Table 7: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Table 8: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Table 9: Alignment rank of LLMs over different age groups in Gemany and Great Britain. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Table 10: Alignment rank of LLMs over different age groups in Indonesia and Malaysia. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.